# Electro-optically tunable second-harmonic generation in lithium niobate metasurfaces boosted by Brillouin zone folding induced bound states in the continuum

### v1

@@ Abstract
Lithium niobate (LN) metasurfaces exhibit remarkable Pockels effect-driven electro-optic tunability, enabling dynamic control of optical responses through external electric fields. When combined with their high second-order nonlinear susceptibility (), this tunability is projected into the nonlinear landscape, realizing second-harmonic generation (SHG)-dominated functionalities in integrated photonics. However, achieving deep SHG modulation in LN metasurfaces remains challenging due to LN’s limited refractive index tunability under practical driving voltage. To address this, we design an air-hole-structured LN metasurface by strategically adjusting air hole positions to induce Brillouin zone folding-enabled bound states in the continuum with ultrahigh quality factors (). Numerical simulations demonstrate a SHG conversion efficiency at 2 MW/cm2 excitation and a modulation depth exceeding 0.99 under 12 V peak-to-peak voltage (). This work establishes a compact framework for electrically tunable nonlinear optics, advancing applications in integrated quantum light sources and programmable photonic chips.
@@ I Introduction
In this work, to achieve efficient modulation of SHG in LN metasurfaces, we synergistically integrate the superior electro-optic tunable nonlinear properties of LN with the unique advantages of BZF-BIC. The metasurface design comprises a unit cell of periodically patterned air holes in a LN thin film, featuring group symmetry. By precisely adjusting the relative positions of air holes to introduce periodic structural perturbations, we modified the metasurface unit cell configuration, thereby inducing BZF-BIC with high -factor () and ultra-narrow resonance linewidths. Numerical simulations demonstrate a SHG conversion efficiency at an input pump intensity of 2 MW/cm2. Furthermore, on-to-off SHG switching (modulation depth 0.99) is achieved under a peak-to-peak modulation voltage of 12 V. This work provides a paradigm for designing nonlinear optical modulators, tunable optical filters, and all-optical switches based on metasurfaces, significantly advancing the development of high-efficiency integrated dynamic nonlinear photonic systems.
@@ II.1 BZF-BIC in LN metasurfaces
The inset of Fig. 2(a) schematically illustrates the BZF process induced by periodic structural perturbations. By displacing the central air hole along the -axis, the M point at the Brillouin zone boundary folds along the M-X’- path toward the point, generating a new type of BIC termed BZF-BIC. To validate this mechanism, we perform finite element simulations in COMSOL Multiphysics with periodic boundary conditions along the and directions (modeling infinite periodicity), and perfectly matched layers (PMLs) in the direction (see Supplemental Material Section 3). Noted, LN’s refractive index is modeled with full dispersion consideration rather than as a constant, essential for capturing electro-optic modulation effects in later sections (see Supplemental Material Section 1). As shown in Fig. 2(a), the single-hole structure exhibits a guided mode at the M point with a eigenfrequency of THz. Upon introducing perturbations (double-hole configuration), this mode’s energy band folds through X’ to , which is consistent with the BZF process—confirming BZF-BIC formation in the inset (see Supplemental Material Section 2). Figure 2(b) demonstrates that the mode’s -factor follows , aligning with symmetry-protected BIC scaling laws (see Supplemental Material Section 4)[60]. Figure 2(c) further reveals the mode’s electromagnetic field symmetry at the point: the electric field components and exhibit odd symmetry, while the magnetic field component shows even symmetry. These features provide conclusive evidence for the formation of the BZF-BIC.
@@ II.2 Nonlinear SHG
We calculate the SHG spectra of the metasurface under a pump intensity of 2 MW/cm2 for different perturbation displacements. As shown in Fig. 3(a), increases as decreases. This trend is closely related to the geometric asymmetry, the -factor of the BZF-BIC, and the enhanced field confinement at smaller . Figure 3(b) explicitly correlates with . Theoretically, as approaches zero, the -factor tends toward infinity, significantly enhancing the local field and enabling to reach extremely high levels. However, in practical fabrication processes, unavoidable manufacturing errors can degrade metasurface performance, necessitating a trade-off between achieving a high -factor and ensuring fabrication feasibility. When nm (a value achievable under current fabrication conditions[62, 63]), the metasurface achieves an SHG efficiency of .
On the other hand, is also influenced by the pump intensity. As shown in Fig. 3(c), exhibits a linear relationship with the pump intensity , indicating that increasing can enhance by several orders of magnitude. This result aligns well with previous theoretical predictions of second-order nonlinear process enhancement based on quasi-BIC[42, 47, 48, 49, 50]. However, if continues to increase, will be affected by backward frequency conversion and thus no longer exhibits linear enhancement (see Supplemental Material Section 6). Furthermore, we investigated the effect of the pump polarization on . As shown in Fig. 3(d), remains high for -polarized pump light but nearly vanishes for -polarized light, demonstrating strong polarization dependence. This phenomenon is primarily due to two factors: First, introducing the perturbation factor breaks the symmetry of the structure, resulting in differences in linear transmission spectra for different polarization directions and consequently leading to polarization-dependent SHG. Second, although the -cut lithium niobate metasurface exhibits an isotropic refractive index in the - plane, the nonlinear polarization terms in the -direction, expressed as , are affected differently by and . This also contributes to polarization-dependent SHG characteristics.
@@ II.3 Electro-optic modulation
Based on the effects of the perturbation displacement , the pump intensity , and the polarization dependence on the nonlinear conversion efficiency of the LN metasurface discussed in the previous sections, in the following electro-optic modulation, we conduct simulations where -polarized light with a pump intensity of 2 MW/cm2 is incident on the LN metasurface with nm. Due to the linear Pockels effect in LN, the refractive index as a function of the applied electric field is given by[64]
@@
Figure 4(a) clearly shows that applying a positive voltage induces a blue shift in the resonance wavelength, while a negative voltage causes a red shift. This behavior stems from the negative linear correlation between LN’s refractive index and the applied voltage. The resonance wavelength shift directly affects the spectral position of the SHG, as demonstrated in Fig. 4(b), which displays SHG spectral shifts under V applied voltage. To visualize the switching modulation effect of the LN metasurface at V, Fig. 4(c) plots the absolute difference between the SHG efficiency at V and V, revealing an ’M’-shaped distribution. After normalizing the data for quantitative analysis, the modulation depth approaches unity, confirming that the on-to-off SHG switching control is achieved with a 12 V voltage difference.
@@
This work | BZF-BIC | 1621 | 12 | YES |
We further examine the LN metasurface’s modulation performance under varying applied voltages. Figure 4(d) plots the modulation depth versus voltage difference in the range of 0–12 V, showing a progressive enhancement of the modulation effect with increasing voltage until saturation occurs at V, where stable switching modulation is attained. Further voltage increases yield negligible improvement. Additionally, the modulation voltage cannot exceed LN’s damage threshold. Within this material-specific limit, electric-field-induced resonance wavelength shifts enable precise tuning of SHG peaks to target wavelengths. Compared to structural tuning via perturbation modifications, this electro-optic approach avoids the cumbersome re-fabrication process while preserving high SHG efficiency, fully demonstrating reconfigurable metasurfaces’ capability for nonlinear optical control.
@@ Acknowledgements.
- Carletti et al. [2019] L. Carletti, C. Li, J. Sautter, I. Staude, C. De Angelis, T. Li, and D. N. Neshev, Second harmonic generation in monolithic lithium niobate metasurfaces, Opt. Express 27, 33391 (2019).

### v2

@@ Abstract
Lithium niobate (LN) metasurfaces exhibit remarkable Pockels effect-driven electro-optic tunability, enabling dynamic control of optical responses through external electric fields. When combined with their high second-order nonlinear susceptibility (), this tunability is projected into the nonlinear landscape, realizing second-harmonic generation (SHG)-dominated functionalities in integrated photonics. However, achieving deep SHG modulation in LN metasurfaces remains challenging due to LN’s limited refractive index tunability under practical driving voltage. To address this, we design an air-hole-structured LN metasurface by strategically adjusting air hole positions to induce Brillouin zone folding-enabled bound states in the continuum with ultrahigh quality factors (). Numerical simulations demonstrate a SHG conversion efficiency at 0.1 MW/cm2 excitation and a modulation depth exceeding 0.99 under 15 V peak-to-peak voltage (). This work establishes a compact framework for electrically tunable nonlinear optics, advancing applications in integrated quantum light sources and programmable photonic chips.
@@ I Introduction
In this work, to achieve efficient modulation of SHG in LN metasurfaces, we synergistically integrate the superior electro-optic tunable nonlinear properties of LN with the unique advantages of BZF-BIC. The metasurface design comprises a unit cell of periodically patterned air holes in a LN thin film, featuring group symmetry. By precisely adjusting the relative positions of air holes to introduce periodic structural perturbations, we modified the metasurface unit cell configuration, thereby inducing BZF-BIC with high -factor () and ultra-narrow resonance linewidths. Numerical simulations demonstrate a SHG conversion efficiency at an input pump intensity of 0.1 MW/cm2. Furthermore, on-to-off SHG switching (modulation depth ) is achieved under a peak-to-peak modulation voltage of 15 V. This work provides a paradigm for designing nonlinear optical modulators, tunable optical filters, and all-optical switches based on metasurfaces, significantly advancing the development of high-efficiency integrated dynamic nonlinear photonic systems.
@@ II.1 BZF-BIC in LN metasurfaces
The inset of Fig. 2(a) schematically illustrates the BZF process induced by periodic structural perturbations. By displacing the central air hole along the -axis, the M point at the Brillouin zone boundary folds along the M-X’- path toward the point, generating a new type of BIC termed BZF-BIC. To validate this mechanism, we perform finite element simulations in COMSOL Multiphysics with periodic boundary conditions along the and directions (modeling infinite periodicity), and perfectly matched layers (PMLs) in the direction (see Supplemental Material Section 3). Noted, LN’s refractive index is modeled with full dispersion consideration rather than as a constant, essential for capturing electro-optic modulation effects in later sections (see Supplemental Material Section 1). As shown in Fig. 2(a), the single-hole structure exhibits a guided mode at the M point with a eigenfrequency of THz. Upon introducing perturbations (double-hole configuration), this mode’s energy band folds through X’ to , which is consistent with the BZF process—confirming BZF-BIC formation in the inset (see Supplemental Material Section 2). Figure 2(b) demonstrates that the mode’s -factor follows the inverse square law relationship, aligning with symmetry-protected BIC scaling laws (see Supplemental Material Section 4)[60]. Figure 2(c) further reveals the mode’s electromagnetic field symmetry at the point: the electric field components and exhibit odd symmetry, while the magnetic field component shows even symmetry. These features provide conclusive evidence for the formation of the BZF-BIC.
@@ II.2 Nonlinear SHG
We calculate the SHG spectra of the metasurface under a pump intensity of 0.1 MW/cm2 for different perturbation displacements. As shown in Fig. 3(a), increases as decreases. This trend is closely related to the geometric asymmetry, the -factor of the BZF-BIC, and the enhanced field confinement at smaller . Figure 3(b) explicitly correlates with . Theoretically, as approaches zero, the -factor tends toward infinity, significantly enhancing the local field and enabling to reach extremely high levels. However, in practical fabrication processes, unavoidable manufacturing errors can degrade metasurface performance, necessitating a trade-off between achieving a high -factor and ensuring fabrication feasibility. When nm (a value achievable under current fabrication conditions[62, 63]), the metasurface achieves an SHG efficiency of .
On the other hand, is also influenced by the pump intensity. As shown in Fig. 3(c), exhibits a linear relationship with the pump intensity , indicating that increasing can enhance by several orders of magnitude. This result aligns well with previous theoretical predictions of second-order nonlinear process enhancement based on quasi-BIC[42, 47, 48, 49, 50]. However, if continues to increase, will be affected by backward frequency conversion and thus no longer exhibits linear enhancement (see Supplemental Material Section 6). Furthermore, we investigated the effect of the pump polarization on . As shown in Fig. 3(d), remains high for -polarized pump light but nearly vanishes for -polarized light, demonstrating strong polarization dependence. This phenomenon is primarily due to two factors: First, introducing the perturbation factor breaks the symmetry of the structure, resulting in differences in linear transmission spectra for different polarization directions and consequently leading to polarization-dependent SHG. Second, in addition to the slight difference in the refrative index in the - plane, the nonlinear polarization terms are affected differently by and . This also contributes to polarization-dependent SHG characteristics.
@@ II.3 Electro-optic modulation
Based on the effects of the perturbation displacement , the pump intensity , and the polarization dependence on the nonlinear conversion efficiency of the LN metasurface discussed in the previous sections, in the following electro-optic modulation, we conduct simulations where -polarized light with a pump intensity of 0.1 MW/cm2 is incident on the LN metasurface with nm. Due to the linear Pockels effect in LN, the refractive index as a function of the applied electric field is given by[64]
@@
Figure 4(a) clearly shows that applying a positive voltage induces a blue shift in the resonance wavelength, while a negative voltage causes a red shift. This behavior stems from the negative linear correlation between LN’s refractive index and the applied voltage. The resonance wavelength shift directly affects the spectral position of the SHG, as demonstrated in Fig. 4(b), which displays SHG spectral shifts under V applied voltage. To visualize the switching modulation effect of the LN metasurface at V, Fig. 4(c) plots the absolute difference between the SHG efficiency at V and V, revealing an ’M’-shaped distribution. After normalizing the data for quantitative analysis, the modulation depth approaches unity, confirming that the on-to-off SHG switching control is achieved with a 15 V voltage difference.
@@
This work | BZF-BIC | 1606 | 15 | YES |
We further examine the LN metasurface’s modulation performance under varying applied voltages. Figure 4(d) plots the modulation depth versus voltage difference in the range of 0–15 V, showing a progressive enhancement of the modulation effect with increasing voltage until saturation occurs at V, where stable switching modulation is attained. Further voltage increases yield negligible improvement. Additionally, the modulation voltage cannot exceed LN’s damage threshold. Within this material-specific limit, electric-field-induced resonance wavelength shifts enable precise tuning of SHG peaks to target wavelengths. Compared to structural tuning via perturbation modifications, this electro-optic approach avoids the cumbersome re-fabrication process while preserving high SHG efficiency, fully demonstrating reconfigurable metasurfaces’ capability for nonlinear optical control.
@@ Acknowledgements.
- Seres [2001] J. Seres, Dispersion of second-order nonlinear optical coefficient, Appl. Phys. B 73, 705 (2001).